### v2

# Diffusion Models for Image Restoration and Enhancement: A Comprehensive Survey

Abstract
Image restoration (IR) has been an indispensable and challenging task in the low-level vision field, which strives to improve the subjective quality of images distorted by various forms of degradation. Recently, the diffusion model has achieved significant advancements in the visual generation of AIGC, thereby raising an intuitive question, “whether the diffusion model can boost image restoration”. To answer this, some pioneering studies attempt to integrate diffusion models into the image restoration task, resulting in superior performances than previous GAN-based methods. Despite that, a comprehensive and enlightening survey on diffusion model-based image restoration remains scarce. In this paper, we are the first to present a comprehensive review of recent diffusion model-based methods on image restoration, encompassing the learning paradigm, conditional strategy, framework design, modeling strategy, and evaluation. Concretely, we first introduce the background of the diffusion model briefly and then present two prevalent workflows that exploit diffusion models in image restoration. Subsequently, we classify and emphasize the innovative designs using diffusion models for both IR and blind/real-world IR, intending to inspire future development. To evaluate existing methods thoroughly, we summarize the commonly used dataset, implementation details, and evaluation metrics. Additionally, we present the objective comparison for open-sourced methods across three tasks, including image super-resolution, deblurring, and inpainting. Ultimately, informed by the limitations in existing works, we propose nine potential and challenging directions for the future research of diffusion model-based IR, including sampling efficiency, model compression, distortion simulation and estimation, distortion invariant learning, and framework design. The repository is released at https://github.com/lixinustc/Awesome-diffusion-model-for-image-processing/
Keywords:
Diffusion models Image restoration Image enhancement Image processing
journal: International Journal of Computer Vision
1 Introduction
Image Restoration (IR) has been a long-term research topic in low-level vision tasks, which plays an irreplaceable role in improving the subjective quality of images. The popular IR tasks contain image super-resolution (SR) (Dong et al., 2014; Lim et al., 2017a; Dai et al., 2019; Wang et al., 2018a; Lu et al., 2022; Chen et al., 2023b; Zhang et al., 2022a; Gao et al., 2023a), deblurring (Nah et al., 2017a; Kupyn et al., 2018; Nimisha et al., 2017; Kim et al., 2023b; Zamir et al., 2022; Tsai et al., 2022), denoising (Zhang et al., 2017; Valsesia et al., 2020; Fan et al., 2022; Liu et al., 2022a; Tian et al., 2023; Zhang et al., 2023a), inpainting (Xie et al., 2012; Pathak et al., 2016; Zhao et al., 2020; Yu et al., 2023b; Li et al., 2022c) and compression artifacts removing (Dong et al., 2015; Svoboda et al., 2016; Ehrlich et al., 2020; Wang et al., 2022b; Jiang et al., 2023c), etc. To restore the distorted images, traditional IR methods treated the restoration as the signal processing and reduce the artifacts with hand-crafted algorithms from the spatial or frequency perspective (Dabov et al., 2007; Chantas et al., 2009; Mairal et al., 2008; Xu and Sun, 2010; Schmidt and Roth, 2014). With the development of deep learning, numerous IR works collected a series of datasets tailored for various IR tasks, e.g., DIV2K (Agustsson and Timofte, 2017), LSDIR (Li et al., 2023d), Set5 (Bevilacqua et al., 2012), and Set14 (Zeyde et al., 2012) for SR, Rain800 (Zhang et al., 2019a), Rain200 (Yang et al., 2017), Raindrop (Qian et al., 2018) and DID-MDN (Zhang and Patel, 2018) for draining, REDS (Nah et al., 2019), and Gopro (Nah et al., 2017b) for motion deblurring, etc. Leveraging these datasets, the majority of recent works (Lim et al., 2017a; Nah et al., 2017a; Valsesia et al., 2020; Dong et al., 2015; Svoboda et al., 2016; Zhang et al., 2017; Lu et al., 2022; Liu et al., 2021; Zhang et al., 2022a; Wang et al., 2022e; Zamir et al., 2022) focused on improving the representation capability of IR networks for complicated degradation through well-designed backbones based on Convolutional neural networks (CNNs) (Krizhevsky et al., 2017) or Transformer (Vaswani et al., 2017). Although these works achieve great progress in the objective quality (e.g., PSNR, and SSIM), the restored images still suffer from unsatisfactory texture generation, hindering the application of these IR methods in real-world scenarios.
Thanks to the development of generative models (Kingma and Welling, 2013; Van Den Oord et al., 2016; Salimans et al., 2017; Rezende and Mohamed, 2015; Papamakarios et al., 2017; Creswell et al., 2018; Karras et al., 2017a), especially the generative adversarial network (GAN) (Creswell et al., 2018), some pioneering IR studies (Dosovitskiy and Brox, 2016; Yu and Porikli, 2016; Johnson et al., 2016; Ledig et al., 2017; Wang et al., 2018a) pinpoint that previous pixel-wise loss, e.g., MSE loss, and L1 loss, are susceptible to the blurring textures, and introduce the adversarial loss from GAN to the optimization of IR network, for enhancing its texture generation capability. For instance, SRGAN (Ledig et al., 2017) and DeblurGAN (Kupyn et al., 2018) utilize a combination of pixel-wise loss and adversarial loss to achieve perception-oriented SR network and deblurring network, respectively. Following them, two primary directions to improve GAN-based IR arise by enhancing the generator (i.e., the restoration network) (Ledig et al., 2017; Zhang et al., 2019b; Wang et al., 2018a, 2021b; Zhang et al., 2021a) and discriminator (Schonfeld et al., 2020; Isola et al., 2017; Karras et al., 2020). In particular, ESRGAN (Wang et al., 2018a) introduces the powerful RRDB (Wang et al., 2018a) as a generator for GAN-based SR tasks. Three popular discriminators, including pixel-wise discriminator (U-Net shape) (Schonfeld et al., 2020), patch-wise discriminator (Isola et al., 2017; Wang et al., 2018b), and image-wise discriminator (Simonyan and Zisserman, 2014; Karras et al., 2020) (i.e., VGG-like architecture) are designed to focus on the subjective quality at different levels of granularity, (i.e., from local to global). Although the above progress, most researches on GAN-based IR still face two inevitable but crucial problems: 1) the training of GAN-based IR is susceptible to mode corruption and unstable optimization and 2) the textures of most generated images seem to be fake and counterfactual.
Recently, diffusion models have emerged as a new branch of generative models, leading to a series of breakthroughs in visual generation tasks. The prototype of the diffusion model can be traced back to the work (Sohl-Dickstein et al., 2015a), and has been developed by DDPM (Ho et al., 2020), NCSN (Song and Ermon, 2019) and SDE (Song et al., 2020b). In general, the diffusion model is composed of the forward/diffusion process and reverse process, where the forward process progressively increases the pixel-wise noise to the image until it satisfies the distribution of Gaussian noise, and the reverse process aims to reconstruct the image by denoising with score estimating (Song and Ermon, 2019) or noise prediction (Ho et al., 2020). Compared to GANs, the diffusion models yield high-fidelity and diverse generation results, thereby successfully replacing the GANs in a series of fields, such as visual generation (Song and Ermon, 2019; Song et al., 2020a, b; Ho et al., 2020) and conditional visual generation (Singh et al., 2022; Popov et al., 2021; Babnik et al., 2023; Blattmann et al., 2023). With the advancement of the vision-language model, diffusion model has been extended to the cross-modality generation, such as StableDiffusion (Rombach et al., 2022), and DALLE-2 (Ramesh et al., 2022). This greatly promotes the development of artificial intelligence generated content (AIGC). We have listed the representative works on diffusion models based on the timeline in Fig. 1 of the Supplementary.
Inspired by the superior generative capability of diffusion models, numerous studies have investigated their application in image restoration tasks, targeting to facilitate the texture recovery. According to the training strategy, these works can be roughly divided into two categories: 1) the first category (Saharia et al., 2022b; Li et al., 2022b; Dos Santos et al., 2022; Özdenizci and Legenstein, 2023; Whang et al., 2022; Luo et al., 2023c; Zhou et al., 2023; Chan et al., 2023; Miao et al., 2023; Varanka et al., 2024; Qu et al., 2024; Zhang et al., 2023d) is dedicated to optimizing the diffusion model for IR from scratch via supervised learning, and 2) the second one (i.e., zero-shot one) (Kawar et al., 2021, 2022b, 2022a; Wang et al., 2023g; Chung et al., 2023b; Song et al., 2021; Liu et al., 2023b; Abu-Hussein et al., 2022) strives to exploit the generative priors in the pre-trained diffusion models for IR. Typically, supervised learning-based methods necessitate the collection of large-scale distorted/clean image pairs, while zero-shot-based methods predominantly rely on known degradation modes. These limitations impede the application of these diffusion model-based methods in real-world scenarios, where the distortions are typically diverse and unknown. To further tackle the above issue, some researches (Sahak et al., 2023; Murata et al., 2023; Chung et al., 2023a; Wang et al., 2023c; Wei et al., 2023; Wang et al., 2023i) have extended the diffusion model to handle blind/real-world image restoration by incorporating real-world distortion simulation, kernel estimation, domain translation, and distortion invariant learning.
Although the diffusion models have shown significant efficacy in image restoration, the associated techniques and benchmarks exhibit considerable diversity and complexity, making them hard to be followed and improved. Moreover, the absence of a comprehensive review for diffusion model-based IR further limits its development. In this paper, we are the first to review and summarize the works on diffusion model-based image restoration methods, aiming to provide a well-structured and in-depth knowledge base and facilitate its evolution within the image restoration community. More analysis on the advantages of our survey can be found in the Section E of our Supplementary.
In this survey, we start by presenting the background of diffusion models in Sec. 2, highlighting two foundational modeling methods, i.e., DDPM (Ho et al., 2020), and SDE (Song et al., 2020b). Based on these preliminaries, we shed light on the advances of diffusion models in image restoration from two distinct directions in Sec. 3: 1) supervised diffusion model-based IR, and 2) zero-shot diffusion model-based IR. In Sec. 4, we summarize the diffusion model-based IR under more practical and challenging scenarios, i.e., blind/real-world degradation. This intends to further enhance the capability of the diffusion model-based IR methods in fulfilling the demands of practical application. To facilitate a reasonable and exhaustive comparison, in Sec. 5, a series of comparisons between different DM-based benchmarks across four tasks are provided, including image SR, inpainting, and deblurring, and real-world restoration. In Sec. 6, we delve into analyzing the primary challenges and potential directions in diffusion model-based IR to provide some insights and hope to inspire more new works. The final conclusion for this review is summarized in Sec. 7.
2 Background on Diffusion Model (DM)
Diffusion probabilistic model (i.e., diffusion model) has brought an evolution in the field of generative models, which transforms the complicated and unstable generation process into several independent and stable reverse processes via Markov Chain modeling. There are three foundational diffusion models that are widely utilized, including DDPM (Ho et al., 2020). NCSNs (Song and Ermon, 2019) and SDE (Song et al., 2020b). Among them, NCSNs (Song and Ermon, 2019) seeks to model the data distribution by sampling using annealed Lange-vin dynamics with a sequence of decreasing noise scales. In contrast, DDPM (Ho et al., 2020) models the forward process with a fixed process of adding Gaussian noise, which simplifies the reverse process of the diffusion model into a solution process for the variational bound objective. These two basic diffusion models are actually special cases of score-based generative models (Song et al., 2020b). SDE (Song et al., 2020b), as the unified form, models the continuous diffusion and reverse processes with stochastic differential equation (SDE). It proves that the NCSNs and DDPM are only two separate discretizations of SDE. We will clarify two commonly-used modeling strategies of diffusion models in image restoration, i.e., DDPM (Ho et al., 2020) and SDE (Song et al., 2020b) in the following subsections. We also illustrate the NCSN (Song and Ermon, 2019) in the section B.1 of the Supplementary.
2.1 Denoising Diffusion Probabilistic Model
DDPM (Denosing diffusion probabilistic model) (Ho et al., 2020) originates from the diffusion models (Sohl-Dickstein et al., 2015b), which introduces the simple variational bound objective for diffusion models by setting the variance $\beta_{t}$ as fixed values. There are two crucial processes in diffusion models, i.e., the forward process and reverse process. In particular, the forward process (i.e., the diffusion process in DDPM) aims to progressively corrupt the training data to the Gaussian Noise, which is a parameterized Markov chain as:
$q(x_{t}|x_{t-1})=\mathcal{N}(x_{t};\sqrt{1-\beta_{t}}\cdot x_{t-1},\beta_{t}\mathbf{I}),$ (1)
where $x_{0},x_{1},...,x_{T}$ are the noise latent variables by adding noises to the training data point $x_{0}\sim p_{data}(x)$ progressively with noise schedule as $\beta_{1},...,\beta_{T}\in(0,1)$ for $T$ steps. And we can compute the probabilistic distribution of $x_{t}$ given $x_{0}$ as:
$q(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\hat{\alpha}_{t}}x_{0},\sqrt{1-\hat{\alpha}_{t}}\mathbf{I}),$ (2)
where $\alpha_{t}=1-\beta_{t}$ and $\hat{\alpha}_{t}=\prod\limits_{i=1}^{t}\alpha_{i}$. When time step $t\to T$ is large enough, the distribution of $x_{T}$ will be a standard Gaussian distribution $\pi(x_{T})\sim\mathcal{N}(0,\mathbf{I})$ since $\hat{\alpha}_{t}\to 0$.
The reverse process of diffusion models aims to recover the data distribution from the Gaussian noises by approximating the posterior distribution $q(x_{t-1}|x_{t},x_{0})$ as:
$q(x_{t-1}|x_{t},x_{0})=\mathcal{N}(x_{t-1};\tilde{\mu}_{t}(x_{t},x_{0}),\tilde{\beta}_{t}\mathbf{I}),$ (3)
where $\tilde{\mu}_{t}(x_{t},x_{0})=\frac{\sqrt{\hat{\alpha}_{t-1}}\beta_{t}}{1-\hat{\alpha}_{t}}x_{0}+\frac{\sqrt{\hat{\alpha}_{t}}(1-\hat{\alpha}_{t-1})}{1-\hat{\alpha}_{t}}x_{t}=\frac{1}{\sqrt{\alpha_{t}}}(x_{t}\\ -\frac{\beta_{t}}{\sqrt{1-\hat{\alpha}_{t}}})\epsilon$ for $\epsilon\sim\mathcal{N}(\mathbf{0},\textbf{I})$ and $\tilde{\beta_{t}}=\frac{1-\hat{\alpha}_{t-1}}{1-\hat{\alpha}_{t}}$. As stated in Eq. 3, the variance schedule $\beta_{t}$ is predefined, and thus, it only requires approximating the mean $\mu_{\theta}(x_{t},t)=\tilde{\mu}_{t}(x_{t},x_{o})$ by a denoising network $\epsilon_{\theta}(x_{t},t)$. The optimization objective (Ho et al., 2020) for denoising network can be written as:
[…]
The direct illustration of denoising diffusion probailisitic models(DDPM) is shown in Fig. 1
Figure 1: Denoising Diffusion Probabilistic Models.
2.2 Stochastic Differential Equations (SDEs)
To unify the approaches with the score-based generative modeling and diffusion probabilistic modeling, SDEs (Song et al., 2020b) exploit the continuous diffusion process through the stochastic differentiable equation (SDE) as:
$dx=\mathbf{f}(x,t)dt+g(t)d\mathbf{w}$ (5)
where $\mathbf{w}$ is the standard Wiener process, $\mathbf{f}(\cdot,t)$ is called drift coefficient of $x(t)$ and $g(\cdot)$ is called the diffusion coefficient of $x(t)$. Here, the diffusion coefficient can be understood as the degree perturbed by random noise, and the drift coefficients can be designed to ensure the Gaussian distribution, such as DDPM (Ho et al., 2020) and NCSN (Song and Ermon, 2019). The reverse process of the above continuous diffusion process (i.e., sampling data from noise) is also a diffusion process and can be modeled by a reverse-time SDE:
$dx=[\mathbf{f}(x,t)-g(t)^{2}\nabla_{x}\log p_{t}(x)]dt+g(t)d\hat{\mathbf{w}}$ (6)
Here $dt$ is an infinitesimal negative time step and $\hat{\mathbf{w}}$ is the standard Wiener process when time flows backward from $T$ to 0. The core of reverse-time SDE is to estimate the score function using a neural network and then Eq. 6 can be solved with the score matching (Hyvärinen, 2007; Hyvärinen and Dayan, 2005).
Figure 2: The overview of diffusion model-based image restoration models. This figure categorizes diffusion models into two types based on their training methods, namely the supervised-based models (indicated with a blue background) and zero-shot-based models (indicated with an orange background). Additionally, the figure provides a more detailed classification of models within these two categories according to how conditions are incorporated. The models with a green background are specifically designed for real-world image restoration. All works in this figure can be found in Section C, Table 1 and Table 2 of the Supplementary).
DDPM and NCSN can be regarded as discretizations of two different SDEs. When the time variable turns to infinity, the forward process of DDPM will converge to the following:
$dx=-\frac{1}{2}(1-\hat{\alpha}_{t})\mathbf{x}dt+\sqrt{1-\hat{\alpha}_{t}}d\mathbf{w}.$ (7)
And the SDE form of NCSN is as follows:
$dx=\sqrt{\frac{d[\sigma^{2}(t)]}{dt}}d\mathbf{w}.$ (8)
Here, Eq. 7 and Eq. 8 are called the Variance Preserving (VP) SDE and the Variance Exploding (VE) SDE, respectively. More background information about the diffusion model is shown in section B of the Supplementary.
3 Diffusion model-based Image Restoration Methods
According to whether the diffusion models (DMs) are training-free for IR, we can preliminarily classify the DM-based IR methods into two categories, i.e., supervised DM-based methods (Saharia et al., 2022b; Whang et al., 2022; Jin et al., 2022; Qiu et al., 2023; Wang et al., 2023c; Zhou et al., 2023; Jiang et al., 2023b; Chen et al., 2023d; Chan et al., 2023; Yi et al., 2023; Gou et al., 2023; Xu et al., 2023; Wang et al., 2023h, 2024d; Zhang et al., 2023c), and zero-shot DM-based methods (Choi et al., 2021; Kawar et al., 2022a; Chung et al., 2023b; Fei et al., 2023; Song et al., 2023; Zhu et al., 2023; Mardani et al., 2023; Fabian et al., 2023; Song et al., 2021). Particularly, the supervised DM-based IR methods entail training the diffusion model from scratch with paired distorted/clean images of IR datasets. Unlike previous GAN-based methods (Yuan et al., 2018; Dinh et al., 2022; Wang et al., 2020b) that directly take distorted images as input, DM-based IR employ the well-designed conditional mechanism to incorporate the distorted images as guidance during the reverse process. Despite its promising texture generation results, this approach encounters two notable limitations: 1) training the diffusion model from scratch relies on a large quantity of paired training data; 2) collecting paired distorted/clean images in the real world is challenging. In contrast, zero-shot DM-based methods offer an appealing alternative, requiring only distorted images and dispensing with the need for retraining diffusion models. Instead of acquiring the restoration capability from the training datasets of IR, it excavates and exploits the structure and texture priors from the pre-trained diffusion models for image restoration. The core idea stems from the intuition that the pre-trained generative models can be viewed as the structure and texture repository, constructed using real-world datasets with large amounts of data, such as ImageNet (Russakovsky et al., 2015b) and FFHQ (Karras et al., 2019). Consequently, an essential challenge faced by zero-shot DM-based IR methods is: how to extract the corresponding perceptual priors while preserving the data structure from distorted images. In the subsequent subsections, we first briefly review the representative supervised DM-based IR method: SR3 (Saharia et al., 2022b), and zero-shot DM-based IR method: ILVR (Choi et al., 2021). Then we further classify these two types of methods from the perspectives of conditional strategy, diffusion modeling, and framework, which are summarized in Table 1 and Table 2 of the Supplementary, respectively. In addition, the overall taxonomy of diffusion models is illustrated in Fig. 2
3.1 SR3 – Representative Supervised DM for IR
Unlike the pure image generation task that synthesizes images from noise, image restoration seeks to generate high-quality images from the corresponding low-quality images. Hence, the pivotal challenge of supervised DM-based IR lies in how to effectively incorporate the degraded/low-quality image into the diffusion model as the condition. Let us denote the degraded image as $y$. The foundational objective of the diffusion models (DMs) for IR is to learn the posterior distribution $p_{\theta}(x_{t-1}|y,x_{t})$ at time step $t$, such that $x_{0}\sim q(x|y)$ and $x$ denotes the corresponding high-quality image. To achieve this, a pioneering supervised DM-based approach, SR3, is introduced with a straightforward condition strategy. Specifically, it directly concatenates the degraded image with the generated image $x_{t}$ at $t$ time step, effectively enabling the conditional image generation for SR.
Figure 3: The Backbone of SR3 network
As depicted in Fig. 3, SR3 follows the typical DDPM (Ho et al., 2020) framework and utilizes the U-Net model as the noise predictor. Given the low-resolution (LR) image $y$, SR3 initially up-samples it to the desired resolution using bicubic interpolation. Subsequently, it concatenates the super-resolved LR image $y$ with the denoised output $x_{t}$ at $t$ time step, serving as the input of the diffusion model to predict the noise for the t1t-1 step. When reaching $t=0$, the diffusion model can deliver an upsampled high-quality image $x_{0}$ of $y$ as $x_{0}\approx x$.
Figure 4: The Backbone of ILVR network
3.2 ILVR – Representative Zero-shot DM for IR
Although the supervised DM-based IR methods have exhibited remarkable performance, the training process entails substantial computational cost and large-scale paired datasets, which might be prohibitive for some researchers. To circumvent this, zero-shot DM-based IR (Choi et al., 2021; Lugmayr et al., 2022; Kawar et al., 2021; Chung et al., 2022b; Zhang et al., 2022c) is proposed to exploit the intrinsic knowledge within the pre-trained diffusion models. Concretely, it is observed that pre-trained diffusion models for image generation, trained with a large number of natural images, encapsulate considerable prior knowledge about rich textures. And thus, these pre-trained diffusion models can be considered as repositories of texture information. The exploration of reusing such prior knowledge for training-free image restoration is an emerging and promising direction in the low-level vision field.
As the initial work, Choi et al.(Choi et al., 2021) introduces the iterative latent variable refinement (i.e., ILVR) method, which leverages the unconditional diffusion model to enable the training-free conditional generation for image SR and image translation. The pivotal innovation of ILVR involves substituting the low-frequency components in denoised output with their counterpart from the reference image. This substitution process, illustrated in Fig. 4, ensures structure and semantic consistency between the generated and reference images, thereby facilitating the conditional generation. Particularly, given a reference image $y$ (e.g., the distorted image in IR), at time step $t$, the ILVR predicts the denoised result for time t1t-1 using the following formula:
$\mathbf{x^{\prime}}_{t-1}=\sigma_{t}\mathbf{z}+\frac{1}{\sqrt{\alpha_{t}}}(\mathbf{x}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\hat{\alpha}_{t}}}\mathbf{\epsilon_{\theta}(\mathbf{x}_{t},t)})$ (9)
where $\epsilon_{\theta}(x_{t},t)$ represents the noise predicted by the denoising network, and $z$ denotes the standard random Gaussian noise that governs the randomness of the generated image. However, this sampling process inevitably produces inconsistent structures/textures in $x_{t}$, which necessitates a refinement to align with the structures/textures in reference image $y$ by low-frequency substitution:
$x_{t-1}=x^{\prime}_{t-1}+\Phi_{N}(y_{t-1})-\Phi_{N}(x^{\prime}_{t-1}).$ (10)
Here, $\Phi_{N}$ denotes the low-pass filter designed to bypass the low-frequency component from input, and $y_{t-1}\sim q(y_{t-1}|y)$ is the diffused state of $y$ at t1t-1 steps. Following ILVR, most zero-shot DM-based IR methods (Chung et al., 2023b; Song et al., 2023; Kawar et al., 2022a; Wang et al., 2023g; Fabian et al., 2023; Feng et al., 2023; Zhang et al., 2023b) predominantly focus on enhancing the refinement strategies within the sampling process, thereby training-free.
3.3 Supervised DM-based IR.
Motivated by SR3 (Saharia et al., 2022b), numerous studies have endeavored to optimize the supervised DM-based IR framework, focusing on enhancing the condition strategy and exploring potential and more efficient generation spaces. With respect to the condition strategy, we categorize these studies into three types based on the conditions: 1) low-quality reference image, 2) pre-processed references, and 3) revising diffusion process. Regarding generation space, the supervised DM-based IR methods can be classified into three distinct groups: image space, residual space, and latent space. If not mentioned, the majority of studies generate the restored images within image space, where the structures and textures are required to be generated directly. In contrast, the residual-space diffusion model focuses on reconstructing the residuals between the low-quality image and its corresponding high-quality image, which simplifies the complexity of generating the whole image. Latent-based methods utilize a well-designed encoder to transform the image into a compact latent space for generation, thereby improving the generation efficiency. This section will elucidate existing studies on supervised DM-based IR in terms of the above three conditional strategies and the last two generation spaces.
3.3.1 Condition with Low-quality Reference Image
As highlighted in Sec. 3.1, incorporating distorted images as conditions is indispensable and crucial for supervised DM-based IR (Chen et al., 2023c; Wang et al., 2023d; Li et al., 2024b). The SR3 method has shown that substantial performance can be achieved through a simple concatenation operation. It utilizes a direct concatenation of low-quality reference images with the denoised result at the t1t-1 step as the condition for noise prediction at the $t$ step. With the same condition strategy, Saharial et al.(Saharia et al., 2022a) propose a unified diffusion model, termed Palette, for image-to-image translation tasks, which achieves excellent performance on image colorization, inpainting, uncropping, and JPEG artifact removal. Furthermore, they investigate the effects of different optimization objectives for sample diversity and highlight the pivotal role of self-attention within U-Net for the diffusion model. Despite their effectiveness, the aforementioned methods are constrained since they only support a fixed resolution for IR once trained. To adapt the diffusion model for the real-world IR with an arbitrary size, Özdenizci et al.(Özdenizci and Legenstein, 2023) divide the degraded image and corresponding sampling results $x_{t}$ into several overlapped patches, and then utilize the patch-wise concatenation as the inputs of the diffusion model for noise prediction. Additionally, to address the inconsistency problem caused by different sampling patches in the overlapped region, this work introduces the mean estimated noise for each pixel within the overlapped region. To enhance the quality of generated images, Ho et al.(Ho et al., 2022) introduce the three-layer cascaded diffusion model based on the SR3 backbone. The first diffusion model is exploited to achieve the class-conditioned low-resolution image generation, and two additional diffusion models are cascaded to super-resolve the low-resolution generated image, resulting in a higher-resolution and more realistic generated image. In addition to the DDPM-based approaches, there is another work (Dos Santos et al., 2022) that explores the various variants of the continuous diffusion model SDE (Song et al., 2020b) for the face super-resolution using predictor-corrector sampling.
Figure 5: The flowchart of DM-based IR methods using pre-processed references, where the low-quality image is first processed by a pre-trained restoration network or trainable preprocessing module. The output references could be features or clean images
3.3.2 Condition with Pre-processed References
Even though direct concatenation with the low-quality image shows promising performance, the artifacts in low-quality images inevitably cause harmful effects on the generation of the diffusion model, especially for severe and diverse distortions. To mitigate this issue, several studies (Yin et al., 2023; Wang et al., 2023e; Shang et al., 2023a; Niu et al., 2023; Huang et al., 2017; Moser et al., 2023a; Guo et al., 2023b; Wang et al., 2024b; Wu et al., 2023) strive to enhance the condition by preprocessing low-quality images with either jointly trained modules or pre-trained restoration networks. As depicted in Fig. 5, these works can be grouped into two categories based on the preprocessing strategies, i.e., condition with pre-processed reference image and feature.
Pre-processed reference image. To alleviate the side effects of artifacts in low-quality images, CDPMSR (Niu et al., 2023) exploits existing super-resolution models, e.g., RCAN (Zhang et al., 2018b), SwinIR (Liu et al., 2021), EDSR (Lim et al., 2017b), to enhance the low-quality image, thereby providing a high-quality and more reliable condition for diffusion model. Additionally, it eschews the stochastic sampling during the reverse process in favor of a deterministic denoising process, yielding superior image quality and faster inference. Notably, the pre-processing reference image can serve as not only an enhanced condition but also an initially well-restored image. Therefore, ResDiff (Shang et al., 2023a) utilizes a pre-trained CNN to generate a low-frequency content-abundant image as the initially restored image and exploits the conditional diffusion model to further generate the residuals between the pre-processed distorted image and its corresponding clean image. In contrast, DehazeDDPM (Yu et al., 2023a) tries to add a physics-aware image as an auxiliary condition to achieve complex dehazing task. Thus they generate the transmission map, and haze-free image $I$ through physical modeling. Then, these preprocessed images are then incorporated into diffusion through dynamic fusion.
Pre-processed reference feature. Another popular line is to use the feature of the reference image as the condition for the diffusion model (Varanka et al., 2024; Guo et al., 2023a; Gao et al., 2023b; Jin et al., 2024; Xiao et al., 2023; Wang et al., 2023e; Liu et al., 2024d). IDM (Gao et al., 2023b), striving for continuous image super-resolution, first extracts the initial features of the low-resolution image with EDSR (Lim et al., 2017b). Then, the initial features are downsampled to multiple scales, which are utilized as the conditions for different upsampling layers in the diffusion model, with the intent of refining the implicit representation. In contrast, ShadowDiffusion (Guo et al., 2023a) leverages a pre-trained transformer backbone to extract the degradation prior (i.e., the degradation-related features) from the distorted reference image. This extracted degradation prior is exploited as the auxiliary to refine the generated shadow mask and serves as the condition for shadow-free image generation.
Figure 6: The architecture of IR-SDE (Luo et al., 2023b), where the forward diffusion process approximately models the image degradation processes. The final output of the forward process is equal to a low-quality image with random noise.
3.3.3 Condition by Revising Diffusion Process
Notably, the above-mentioned supervised DM-based IR methods introduce the conditions by modifying the network while preserving the diffusion process from DDPM (Ho et al., 2020). However, this requires the generation process to start from the noise. To obviate this, some studies (Luo et al., 2023b, c, 2024; Yang et al., 2023a; Zhao et al., 2023b; Wang et al., 2024c, a; Cao et al., 2022; Delbracio and Milanfar, 2023; Xiao et al., 2023; Mirza et al., 2023) condition the diffusion model by modifying the diffusion process, such that the diffused output $x_{T}$ (i.e., the start point during the reverse process) approximates the low-quality image corrupted with few Gaussian noises. As shown in Fig. 6, Luo et al.(Luo et al., 2023b) modify the forward process with mean-reverting SDE to model the unified degradation process of IR as:
$dx=\theta_{t}(\mu-x)dt+\sigma_{t}dw,$ (11)
where $\theta_{t}$ and $\sigma_{t}$ are the time-dependent parameters. $\mu$ denotes the distorted image and $x$ represents its corresponding high-quality image. With the mean-reverting SDE, this work successfully models the image degradation and restoration processes with the modified forward and reverse processes, respectively. This avoids the generation from pure noise and achieves better resto-ration performance. Building on IR-SDE (Luo et al., 2023b), the same team introduces the Refusion (Luo et al., 2023c), which further refines IR-SDE (Luo et al., 2023b) by optimizing aspects such as network architecture, noise level, and denoising steps, etc. To reduce the computation cost, Refusion introduces the U-Net compression strategy, thereby enabling efficient sampling in the latent space.
With the same purpose, Xie et al (Xie et al., 2023) redefine the diffusion process such that the sampling starts from the noisy image. Taking into account the diversity of noise, they derive three separate diffusion processes for the removal of Gaussian, Gamma, and Poisson noises, respectively. In contrast, InDI (Delbracio and Milanfar, 2023) introduces the continuous diffusion process as:
$x_{t}=(1-t)x+ty+t\epsilon n,t\in[0,1],$ (12)
where $x$ and $y$ are the high-quality images and their corresponding low-quality counterparts. $\epsilon$ is a small constant. It can be interpreted as the step-wise interpolation of high-quality and low-quality images at time step $t$, which decomposes the original single-step prediction of supervised image restoration into several small steps, effectively circumventing the regression-to-the-mean effects often found in conventional supervised image restoration. Different from the diffusion process in the spatial domain, HFS-SDE (Cao et al., 2022) reformulate the diffusion process for magnetic resonance (MR) reconstruction at frequency space. In this approach, the forward process progressively adds noise into high-frequency space, resulting in the final $x_{T}$ that is composed of high-frequency noise and low-frequency data. During the reverse process, HFS-SDE employs Predictor-Corrector (PC) method (Song et al., 2020b) for sampling.
Figure 7: The architecture of Predict and Refine model (Whang et al., 2022). The predictor $g_{\theta}$ generates the initial prediction of a clean image, while residual information is modeled by diffusion process.
3.3.4 Generate Residuals
In the supervised DM-based IR, most studies directly generate high-quality images from noise, which necessitates the concurrent generation of structure and textures. However, regenerating the structures/textures already existing in low-quality images unnecessarily burdens the diffusion model and increases extra resource costs. Motivated by this, some representative studies (Li et al., 2022b; Whang et al., 2022; Shang et al., 2023a; Liu et al., 2024b; Yue et al., 2023; Gao and Shan, 2022; Karchev et al., 2022; Liu et al., 2023a) seek to move the generation process from the image space to the residual space. The objective is to generate the residuals between paired high-quality and low-quality images. As the pioneering work, SRDiff (Li et al., 2022b) is the first to utilize the diffusion model to predict the residual in SR. Whang (Whang et al., 2022) introduces a predict-and-refine strategy for image deblurring tasks. As shown in Fig. 7, this work first predicts an initial deblurring image with the deterministic deblurring network and then generates the residual through a stochastic diffusion model. ResDiff (Shang et al., 2023a) mentioned above also adopts this strategy for SR. Resshift (Yue et al., 2023) introduces a novel Markov chain that shifts residuals between high-quality and low-quality images. This residual shifting technique achieves comparable performance even with a reduced number of sampling steps, as few as 15. Another work RDDM (Liu et al., 2024b) redefines the forward diffusion process and incorporates residuals to represent directional diffusion from the target domain to the input domain. RDDM predicts both noise and residuals during training to effectively restore images.
3.3.5 Diffusion on Latent Space
To alleviate the training and sampling costs of the diffusion model, some works (Rombach et al., 2022; Luo et al., 2023c; Xia et al., 2023; Yang et al., 2024a; Chung et al., 2023c; Qu et al., 2024; Luo et al., 2023a; Sun et al., 2023), try to conduct diffuion in latent space. Among them, stableDiffusion (Rombach et al., 2022) is the first work to implement the DM-based generation in latent space. Particularly, it pre-trains an autoencoding model (i.e., an encoder-decoder architecture) to learn the perceptual latent space, which is able to preserve the perceptual quality of the reconstructed image while reducing the computational complexity. Utilizing the pre-trained autoencoder, StableDiffusion transforms the image-wise diffusion process to the latent space, and then introduces various conditions (e.g., text, segmentation map, and image) into the diffusion model with the cross-attention mechanism. Inspired by this, Refusion (Luo et al., 2023c) introduces the latent-wise diffusion model for image restoration to accelerate the training and sampling, which is shown in Fig. 8. In contrast to the above works, where latent space is obtained by compressing the original image, DiffIR (Xia et al., 2023) exploits the latent-wise diffusion model to generate the compact IR priors, which guides dynamic transformer-based restoration network (DIRformer) to achieve better restoration.
Figure 8: An overview of diffusion model used in Refusion (Luo et al., 2023c), the low-quality image is first compressed into latent by an encoder. Then the latent information is modeled by a diffusion process. Here, $F$ denotes the forward process, which progressively adds noise to $x_{0}$, transforming it into $x_{T}$.
3.3.6 Efficient Transfer-Learning
In supervised DM-based IR, efficient and rapid fine-tuning has become a focal point of current research (Wang et al., 2023c; Lin et al., 2024; Yang et al., 2024a; Yu et al., 2024; Jiang et al., 2024b; Ai et al., 2024). This method principally leverages pre-trained models such as DDPM or Stable Diffusion, extracting certain distortion information through feature extraction and subsequently modulating it into the diffusion model. Pioneering work is StableSR, where a time-aware encoder is trained to extract information from low-quality images and incorporate the learned features into pretrained stable diffusion(trained on LAION-5B (Schuhmann et al., 2022)). On the other hand, DiffBIR (Lin et al., 2024) and PASD (Yang et al., 2024b) introduce conditions by training a UNet encoder using zero convolutions, which bears similarities to ControlNet (Zhang and Agrawala, 2023). Another significant advancement in this area is SUPIR (Yu et al., 2024), which builds upon a pre-trained model based on SDXL (Podell et al., 2023). It extracts textual prompts through LLaVA (Liu et al., 2024a) and introduces distorted image information using a trimmed Controlnet.
3.4 Zero-shot DM-based IR.
Different from supervised DM-based IR, zero-shot DM-based IR strives to achieve training-free and data-free image restoration. It generally relies on the pre-trained diffusion models designed for generation tasks and incorporates the condition of low-quality images in the sampling process. The core challenges of this task stem from i) how to maintain the data consistency between low-quality images and generated images, since the pre-trained diffusion models are devoted to preserving the data distribution instead of pixel-wise data consistency; ii) how to excavate the perceptual knowledge aligned with low-quality images, which imposes higher requirements on the design of the condition. In this paper, we roughly summarized the zero-shot DM-based IR methods into three categories, i.e., projection, decomposition, and posterior estimation based methods.
3.4.1 Projection-based Methods.
To mitigate the primary challenges in zero-shot DM-based IR, the projection-based method has been introduced in some studies (Lugmayr et al., 2022; Choi et al., 2021; Chung et al., 2022b; Moser et al., 2023b). This approach aims to extract inherent structures/textures from low-quality images as complementary to generated images at each step, which can ensure data consistency. For instance, the task of image inpainting involves merely generating the content for the mask region. The unmasked region of the low-quality image can substitute the corresponding part of the denoised image at the t1t-1 step, thereby establishing the condition for data consistency during the sampling process. In line with this, RePaint (Lugmayr et al., 2022) exploits a simple projection for image inpainting task:
$x_{t-1}=m\bigodot x_{t-1}^{known}+(1-m)\bigodot x_{t-1}^{unknown},$ (13)
where $x_{t-1}^{known}\sim\mathcal{N}(\sqrt{\hat{\alpha}_{t}}y,1-\hat{\alpha}_{t}\mathbf{I})$ is the diffused results at time step t1t-1 by adding noise for masked image $y$. $x_{t-1}^{unknown}$ is sampled from denoised prediction of diffusion model. In contrast, ILVR (Choi et al., 2021) employs the low-frequency projection for image super-resolution. Theoretically, at time step t1t-1, the predicted latent variable $x_{t-1}$ and $y_{t-1}$ (i.e., adding noise to low-resolution image $y$ at the t1t-1 step in the diffusion process) should share the same low-frequency components. Consequently, it substitutes the low-frequency components of $x_{t-1}$ with its counterpart from $y_{t-1}$, ensuring data consistency and establishing an improved condition for the diffusion model. As an advanced solution, CCDF (Chung et al., 2022b) introduces a unified projection method as:
$x_{t-1}=Ax^{\prime}_{t-1}+b,$ (14)
where $A$ and $b$ are set to achieve data consistency. For instance, in the SR task, the above projection can be instantiated as:
$x_{t-1}=(\mathbf{I}-\mathbf{P})x^{\prime}_{t-1}+y_{t-1},$ (15)
where $\mathbf{P}$ is the degradation matrix of downsampling. Furthermore, this work proves that generation starting from better initialization can boost the speed in the reverse process.
3.4.2 Decomposition-based methods.
It is noteworthy that most image restoration problems can be regarded as linear reverse problems, which can be posed as:
$y=Hx+z,$ (16)
where $H$ is the linear degradation operator and $z$ is a contaminating noise. In this setting, the condition probability $p(x|y)$ cannot be directly estimated since the existence of noise $z$. To get rid of the noise $z$, SNIPS (Kawar et al., 2021) and DDRM (Kawar et al., 2022a) run the diffusion process in the spectral domain with singular value decomposition (SVD) on the degradation operator $H$. In particular, SNIPS (Kawar et al., 2021) is based on the annealed Langevian dynamics and derives the conditional score function on the spectral space, which achieves great performance on image deblurring, super-resolution, and compressive sensing tasks. Following SNIPS, DDRM (Kawar et al., 2022a) further extends the SVD decomposition to the variational objective of linear reverse problems, which reveals a pre-trained DDPM (Ho et al., 2020)/DDIM (Song et al., 2020a) can be the optimal solution for it. Notably, the above works only focus on the linear reverse problem. In contrast, Kawar et al (Kawar et al., 2022b) investigate non-linear inverse problems based on the special case of DDRM (i.e., no noise $z$ in the reverse problem), and extend the pseudo-inverse concept to achieve JPEG artifact correction. For the MRI reconstruction, SVD decomposition is not suitable. To overcome this, Song et al. train the unconditional generative model for medical images from scratch and then exploit the matrix decomposition in the sampling process to solve the linear reverse problem, which is general for unknown measurement processes.
With a different purpose, another decomposition strategy, range-null space decomposition is introduced by DDNM (Wang et al., 2023g) to further improve the zero-shot image restoration, where the range space is responsible for the data consistency, and null space is used to improve the reality (i.e., the perceptual quality). Given the noiseless inverse $y=Hx$, it can be decomposed into:
$y=HH^{\dagger}Hx+H(I-H^{\dagger}H)x,$ (17)
where $H^{\dagger}$ is the pseudo-inverse of degradation operation $H$. We can see the range space $HH^{\dagger}Hx=Hx=y$ can ensure data consistency, and null space $H(I-H^{\dagger}H)x$ has no effects on data consistency since $H(I-H^{\dagger}H)x=\mathbf{0}$. As shown in Fig. 9, based on this, DDNM (Wang et al., 2023g) rectifies the prediction of $x_{0}$ at time step $t$ as: $\hat{x}_{0|t}=H^{\dagger}y+(I-H^{\dagger}H)x_{0|t}$, where $x_{0|t}$ can be estimated with the noise prediction at time step $t$. With the rectified $\hat{x}_{0|t}$, we can compute the denoised output $x_{t-1}$ at time step t1t-1, which ensures the data consistency and serves as a better condition for the next noise prediction. Moreover, DDNM also exploits the SVD to solve the linear inverse problem with noise, termed DDNM+.
Figure 9: The architecture of DDNM (Wang et al., 2023g),the forward degradation operator AA is decomposed into range-null space to meet the realness and data consistency.
3.4.3 Posterior estimation.
The projection-based methods have shown superior performance in the inverse problem of image restoration, where a projection-based measurement consistency correction is added after the reverse sampling step of the diffusion model. However, most projection-based works are devoted to the noiseless inverse problems and usually suffer from unsatisfied data consistency, since the projection throws the sample path off the data manifold (Chung et al., 2022a). To solve the general noisy linear reverse problem, some works (Chung et al., 2022a, 2023b; Fabian et al., 2023; Song et al., 2023; Zhang et al., 2023b; Fei et al., 2023; Cardoso et al., 2023; Garber and Tirer, 2024; Lee et al., 2024; Cao et al., 2024) aim to estimate the posterior distribution $p(x|y)$ with the unconditional diffusion model based on the Bayes theorem. It is equivalent to estimating the conditional posterior $p(x_{t}|y)$ at each step of the reverse process. Based on the Bayes theorem, it can be derived as:
$p(x_{t}|y)=p(y|x_{t})p(x_{t})/p(y).$ (18)
And the corresponding score function can be estimated as:
$\nabla_{x_{t}}\log p_{t}(x_{t}|y)=\nabla_{x_{t}}\log p_{t}(y|x_{t})+s_{\theta}(x,t),$ (19)
where the $s_{\theta}(x,t)$ could be extracted from pre-trained model while the term $p_{t}(y|x_{t})$ is intractable. From the above equation, we can find that the key factor to achieve better solvers for the inverse problem of image restoration is to accurately estimate the $p(y|x_{t})$.
As the pioneering works, MCG (Chung et al., 2022a) and DPS (Chung et al., 2023b) approximate the posterior $p(y|x_{t})$ with $p(y|\hat{x}_{o})$, and $\hat{x}_{o}$ is the expectation given $x_{t}$ as $\hat{x}_{o}=E[x_{o}|x_{t}]$ (Chung et al., 2023b) with Tweedie’s formula. Concretely, MCG (Chung et al., 2022a) considers the data consistency from the perspective of the data manifold, where the manifold constrained gradient is proposed to let the correction lie on the data manifold. However, DPS (Chung et al., 2023b) pinpoints that the projection operation in MCG is harmful to data consistency since it might cause the sampling path off the data manifold. Based on this, DPS (Chung et al., 2023b) discards the projection step in the reverse process and estimates the posterior as:
$\nabla_{x_{t}}\log p_{t}(y|x_{t})$ $\approx\nabla_{x_{t}}\log p(y|\hat{x}_{0})$
$\approx-\frac{1}{\sigma^{2}}\nabla_{x_{t}}\|y-H(\hat{x}_{0}(x_{t}))\|_{2}^{2}$ (20)
[…]
$\nabla_{x_{t}}\log p_{t}(y|x_{t})\approx r_{t}^{-2}((h^{\dagger}(y)-h^{\dagger}(h(\hat{x}_{o})))^{\mathrm{T}}\frac{\partial\hat{x}_{o}}{\partial x_{t}})^{\mathrm{T}}$ (21)
[…]
Different from the above works, some works (Fei et al., 2023; Zhang et al., 2023b) attempt to model $p(y|x_{t})$ with other strategies. It is noteworthy that the higher conditional probability $p(y|x_{t})$ is equivalent to a smaller distance between $D(x_{t})$ and $y$ (Fei et al., 2023). Therefore, GDP (Fei et al., 2023) proposes a heuristic approximation of distribution $p(y|x_{t})$ as follows:
$p(y|x_{t})\approx\frac{1}{Z}\exp(-[s\mathcal{L}(\mathcal{D}(x_{t}),y)+\lambda\mathcal{Q}(x_{t})]),$ (22)
where $\mathcal{L}$ and $\mathcal{Q}$ denote the distance metric and the quality loss, respectively. $Z$ is a normalization factor and $s$ is the scaling factor that controls the weight of guidance. However, the distance $\mathcal{L}$ is hard to be defined since the noise magnitudes in $x_{t}$ and $y$ are different. Thus, they substitute $x_{t}$ with its clean estimation $\hat{x}_{0}$ in the distance measurement. With the same purpose, Copaint (Zhang et al., 2023b) tries to predict the $\hat{x}_{0}$ through the one-step estimation by a neural network.
Instead of modeling intractable distribution $p(y|x_{t})$, Feng et al.(Feng et al., 2023) directly estimates the posterior $p(x_{t}|y)$ from a variational perspective. Following DPI (Sun and Bouman, 2021; Sun et al., 2022), they define a family of distributions $q_{\theta}$ through RealNVP (Dinh et al., 2016) normalizing flow with parameter $\theta$, which is optimized through a minimal KL-divergence between true posterior and estimated distribution $q_{\theta}$.
[…]
4 Diffusion Models for Blind/Real-world Image Restoration
Although the methods in Sec. 3 have achieved great breakthroughs in image restoration, most of them (Chung et al., 2023b; Saharia et al., 2022b; Kawar et al., 2022a; Li et al., 2022b; Saharia et al., 2022a; Özdenizci and Legenstein, 2023; Ho et al., 2022; Song et al., 2023; Wang et al., 2023g) focus on solving the synthetic distortions, which usually perform poorly in the out-of-distribution (OOD) real-world/blind degradations. The reasons stem from the inherent challenges of real-world IR: 1) the unknown degradation modes are hard to be identified. 2) collecting distorted/clean image pairs is non-trivial and even unavailable in the real world. To overcome this, previous works (Wang et al., 2022c; Liang et al., 2022; Wang et al., 2022d) have attempted to solve it by simulating real-world degradations (Wang et al., 2021b; Zhang et al., 2021b, 2022b; Wang et al., 2022c), and unsupervised learning (Wang et al., 2021a; Lugmayr et al., 2019a; Wang et al., 2022d), etc. Inspired by these, some pioneering works (Sahak et al., 2023; Chung et al., 2023a; Wang et al., 2023i; Yin et al., 2022; Abu-Hussein et al., 2022) begin to explore how to exploit diffusion models to solve real-world degradations. In this paper, we divided the DM-based blind/real-world IR (Sahak et al., 2023; Yin et al., 2022; Ren et al., 2022; Wang et al., 2023c; Welker et al., 2022; Yang et al., 2023b; Chan et al., 2023; Wang et al., 2023i; Murata et al., 2023; Chung et al., 2023a; Miao et al., 2023) into four categories, i.e., distortion simulation (Sahak et al., 2023; Yang et al., 2023b), kernel estimation (Chung et al., 2023a; Murata et al., 2023), domain translation (Wei et al., 2023; Yang et al., 2023b), and Distortion-invariant diffusion model (Wang et al., 2023i; Yue and Loy, 2024; Ren et al., 2022).
4.1 Distortion Simulation
Notably, the real-world distortions are usually blind/u-nknown, where the distributions are different from the simple synthetic distortions. For supervised-learning-based IR, this requires the restoration network to have strong generalization capability or the synthetic datasets can cover the real-world distortions. From the causality perspective (Li et al., 2023b), these two purposes all rely on simulating diverse distortions that are similar to real-world distortions, which we called distortion simulation/augmentation. There are several representative DM-based IR methods (Sahak et al., 2023; Yin et al., 2022) utilizing the distortion simulations to improve the robustness of their methods for real-world degradations. The representative method is SR3+ (Sahak et al., 2023), which is based on the diffusion model from SR3 (Saharia et al., 2022b) and introduces the second-order degradations simulation of RealESRGAN (Wang et al., 2021b) for the training. Similarly, to simulate the real-world degradation, Yang et al.(Yang et al., 2023b) propose to synthesize the real-world distorted/clean training pairs using diffusion models, where the distorted images are initialized with second-order degradation in RealESRGAN (Wang et al., 2021b).
4.2 Kernel Estimation
Kernel Estimation is first proposed in blind image restoration (Tzikas et al., 2009; Huang and Xia, 2020; Vasu et al., 2018; Oliveira et al., 2013), where the degradation can be modeled as $y=(x*k)\downarrow_{s}+n$. Here, $k$ is the degradation kernel, and $n$ is the additive noise. Under this setting, the kernel $k$ can be estimated as the guidance to boost the adaptability of the restoration network. Inspired by this, BlindDPS (Chung et al., 2023a) and GibbsDDRM (Murata et al., 2023) attempt to solve the blind inverse problem by estimating the unknown degradation kernel in the sampling process. In particular, BlindDPS (Chung et al., 2023a) exploit the DPS (Chung et al., 2023b) architecture and exploit one parallel diffusion model for the degradation kernel estimation. The diffusion model for kernel estimation is pre-trained on synthetic kernels. Unlike BlindDPS, GibbsDDRM (Murata et al., 2023) achieves the sampling process with partially collapsed Gibbs sampler (Van Dyk and Park, 2008), which samples both kernel parameter and image together from the joint posterior $p(x_{t}|k,y)$.
4.3 Domain Translation
In the real world, it is hard to collect distorted/clean image pairs. Although some works attempt to simulate the degradation process of real distorted images, the distribution of the synthesized distortion is still far away from the real-world one. To further solve the Real-world IR problem, a series of works explore the domain translation techniques for image restoration. Domain Translation (Anoosheh et al., 2018; Lin et al., 2019; Chu and Wang, 2018; Saunders, 2022; Lugmayr et al., 2019b) aims to translate the image from one domain to another domain. From the domain translation perspective, synthetic distorted images, real-world distorted images, and high-quality images can be regarded as three different domains, that share the same contents.
The domain translation-based works for DM-based IR can be roughly divided into two categories: 1) The first line (Yang et al., 2023b) aims to simulate more reliable real-world distorted/clean image pairs by translating the low-quality images from the synthetic domain to the real-world domain. In this way, the simulated datasets can enable the restoration network with better restoration capability for real-world degradation. For instance, Yang et al.(Yang et al., 2023b) is the first to exploit the pre-trained diffusion model to synthesize real-world training pairs, where the diffusion model is pre-trained with real-world low-quality images, and the translation is achieved by warping the synthetic low-quality image to noise space (i.e., the generation inversion). 2) Another popular line (Wei et al., 2023) exploits unsupervised learning, where domain translation is achieved by the cycle consistency constraint. In particular, two generators construct one cycle path, where one generator aims to translate the distorted image to the distortion-free image, and another generator is utilized to translate the clean image to the distorted image. This enables unsupervised training with unpaired real-world distorted and high-quality images. As the representative work, RainDiffusion (Wei et al., 2023) proposes to remove the rain with two cooperative branches, where the non-diffusive translation branch aims to utilize the pre-trained cycle-consistent generators to produce initial paired clean/rainy images, and diffusive translation branch leverages the multi-scale diffusion models to refine the results.
4.4 Distortion-invariant Diffusion Model
Since the blind distortions are usually diverse and complicated, the diffusion model is required to own the generation capability for these distortions (i.e., the distortion-invariant capability) in the real world. To achieve the distortion-invariant diffusion model, DifFace et al.(Yue and Loy, 2024) introduces a pre-trained restoration network, e.g., SRCNN (Dong et al., 2014) or SwinIR (Liu et al., 2021), to obtain an initial clean image as a sampling start point $x_{N}$, where the restoration network is trained with the second-order degradation from RealE-SRGAN (Wang et al., 2021b), thereby exhibiting good generalization capability and produce the distortion invariant initial clean image for diffusion model. Ren et al.(Ren et al., 2022) proposes to achieve a distortion-invariant diffusion model with multi-scale degradation-invariant guidance information. They employ distortion augmentation strategies on degraded images to obtain invariant representation with structure information as guidance. In contrast, Wang (Wang et al., 2023i) exploits the low-pass filter to filter the distortion invariant components in the low-quality image since the different real-world distorted images usually share the same structure information. As shown in Fig. 11, they adopt a simple iterative refinement similar to ILVR during the sampling stage. After obtaining the degradation-invariant $\hat{x}_{0}$, they use an enhancement module (Powerful CNN-based or transformer-based restoration methods) to further improve the image quality.
Figure 11: The architecture of DR2 (Wang et al., 2023i). It consists degradation removal module and an enhancement module. The input image is first sent to a diffusion process to remove degradation information gradually. The output degradation-free image will be enhanced by a pre-trained face restoration network.
4.5 Discussion on Some Real-world Application and Limitation.
Low-light enhancement. Real-world low-light image restoration aims to enhance and restore images captured under poor lighting conditions, which is susceptible to low brightness, high noise, color distortion, and reduced contrast, posing significant challenges for restoration methods. To achieve this, LLDiffusion (Wang et al., 2023e) introduces a degradation generation model to capture the degradation representation for low-light enhancement. This degradation-aware condition enhances the performance of diffusion-based low-light image restoration by providing more informative priors. To improve the efficiency of the diffusion model while preserving the generation capability, DiffLL (Jiang et al., 2023a) decouples the low-light image with wavelet-transform, using the diffusion model only to the low-frequency component. This approach significantly reduces the inference time while preserving the high-frequency details through a dedicated high-frequency restoration module(HFRM). PyDiffusion (Zhou et al., 2023) incorporates a multi-scale generation framework, applying the diffusion model at different scales to mitigate the computational inefficiency of direct high-resolution generation. This hierarchical approach enhances both efficiency and performance. There are also two diffusion-based works incorporating physically inspired priors to serve as conditioning inputs to the diffusion model, improving restoration quality by explicitly modeling the underlying lighting and reflectance properties. For instance, Diff-Retinex (Yi et al., 2023) decomposes low-light images into illumination and reflectance maps based on the Retinex model. LightenDiffusion (Jiang et al., 2024a) introduces an unsupervised training mechanism, it decouples the unpaired normal-light image and low-light images to extract the illumination prior from normal-light image and content reflection from low-light image, then combine them with diffusion model to refine it.
As shown in Table 1, we have compared recent diffusion-based works on real-world low-light image restoration on commonly-used synthetic dataset LOL (Wei et al., 2018) and three commonly-used real-world datasets, including DICM (Lee et al., 2013), NPE (Wang et al., 2013), and VV (Vonikakis et al., 2018). We can observe that DiffLL(Jiang et al., 2023a) and LightenDiffusion(Jiang et al., 2024a) achieve optimal generalization in terms of subjective metrics, i.e., NIQE and PI. However, due to its unsupervised training mechanism, LightenDiffusion exhibits lower objective quality compared to DiffLL (Jiang et al., 2023a).
[…]
Models LOL (Wei et al., 2018) DICM (Lee et al., 2013) NPE (Wang et al., 2013) VV (Vonikakis et al., 2018)
PSNR SSIM LPIPS NIQE PI NIQE PI NIQE PI
CLE Diffusion (Yin et al., 2023) 25.51 0.890 0.160 4.505 3.361 5.249 3.512 3.240 3.470
GDP (Fei et al., 2023) 15.90 0.542 0.337 4.358 3.552 4.032 3.097 4.683 3.431
DiffLL (Jiang et al., 2023a) 26.34 0.845 0.217 3.636 2.936 3.716 2.629 2.351 2.869
LLDiffusion (Wang et al., 2023e) 24.65 0.843 0.075 - - - - - -
Diff-Retinex (Yi et al., 2023) 21.98 0.863 0.048 4.361 3.394 4.996 3.392 3.087 3.350
PyDiffusion (Zhou et al., 2023) 23.28 0.859 0.108 4.499 3.792 4.082 3.268 4.360 3.678
LightenDiffusion (Jiang et al., 2024a) 20.45 0.803 0.192 3.724 3.144 3.618 2.879 2.941 2.558
Adverse weather removal. As the pioneering work, WeathDiff (Özdenizci and Legenstein, 2023) shares the same diffusion architecture with SR3, but it can support arbitrary image restoration by introducing the mean estimated noise of overlapped patches in the reverse process. In contrast, T3-Diffusion (Chen et al., 2024b) defined a group of degradation-specific sub-prompts and utilizes the dynamical composition of them based on unseen degradation for weather removing, which outperforms DA-CLIP (Luo et al., 2024) and Weather-Diff. ReviveDiff (Huang et al., 2024) enhances the perception levels for various weather degradation by introducing the coarse and fine branches into the diffusion model. RainDiff/RainDiffusion (Shen et al., 2023) proposes an unpaired cycle consistency framework for image deraining, eliminating the need for paired training data, making it highly suitable for real-world applications. Depart from that, there are also two diffusion-based works for image dehazing. Dehaze-DDPM (Yu et al., 2023a) employs a two-stage dehazing pipeline, combining the Atmospheric Scattering Model (ASM) for physical modeling in the first stage and diffusion-based generative recovery of lost structural details in the second. Additionally, Wang et al. (Wang et al., 2023b) enhance dehazing by emphasizing mid-to-high-frequency components in the diffusion process, leveraging a frequency-spectrum filter and skip connections to inject high-frequency details into the U-Net architecture.
Shadow removal. ShadowDiffusion (Guo et al., 2023a) leverages a pre-trained transformer backbone to extract the degradation prior (i.e., the degradation-related features) from the distorted reference image. This extracted degradation prior is exploited as the auxiliary to refine the generated shadow mask and serves as the condition for shadow-free image generation. DeS3 (Jin et al., 2024) introduces a novel adaptive attention mechanism to distinguish the underlying objects and shadow regions, which is mask-free compared with existing works on shadow removal. Moreover, it utilizes a pre-trained ViT backbone as loss to enhance the scene understanding capability. As described in DeS3, it can obtain a better performance compared with ShadowDiffusion.
Potential limitations. The potential limitations of current diffusion-based methods in addressing real-world IR are as: (i) Limited task relevance. Some diffusion-based image restoration (IR) methods overlook the inherent characteristics of the tasks themselves and instead rely solely on dataset-specific fitting; (ii) Limited comparison reasonability. As discussed earlier, diffusion-based methods (Özdenizci and Legenstein, 2023; Wei et al., 2023) often use different datasets in some IR task, making it difficult to identify which approach is truly more effective. Furthermore, real-world datasets are challenging to collect and are often insufficient for comprehensive evaluation; (iii) Limited restoration efficiency. As stated in Section 6 of our manuscript, the impressive perceptual quality of diffusion-based IR methods is largely dependent on multiple sampling steps. Although one-step diffusion-based IR methods (Wang et al., 2024c; Wu et al., 2024a) have been proposed, accelerating the sampling process while maintaining restoration quality remains a significant challenge; (iv) Limited pixel-wise consistency. Diffusion-based real image restoration methods (Yu et al., 2024; Jiang et al., 2024a) share a common weakness with GAN-based approaches, i.e., they are susceptible to hallucinated textures due to the generative priors stored within the diffusion model. Addressing this issue remains a critical challenge.
5 Experiments
To ensure the efficient and thorough comparison of different diffusion model-based IR methods, we first summarize the popular datasets, experimental configurations, and evaluation metrics for different tasks. Then we have a comparison of existing benchmarks in several typical image restoration tasks, including image super-resolution, inpainting, deblurring, and JPEG artifacts removal.
5.1 Datasets and Implementation Details
Datasets. It is noteworthy that the contents and degradation modes are significantly different across the datasets from different IR tasks. Therefore, we summarize the commonly-used datasets based on IR tasks, including SR, image deblurring, image inpainting, shadow removal, desnowing, draining, and dehazing in Table 3 of section D of the Supplementary.
Implementation Details. We summarize the implementation details and datasets of supervised algorithms and zero-shot algorithms in Table 4 and Table 5 of the Supplementary, respectively. For supervised algorithms, we describe the configurations in the training process and testing process, including batch size, training iterations, learning rate, the sampling steps in the training process, and the sampling steps in the inference process. For zero-shot-based methods, we clarify the pre-trained diffusion model, evaluated datasets, and sampling steps in the inference process. The commonly used data augmentation strategies are composed of rotation and flip operations.
Table 2: Quantitative Results of Supervised Models on Super-resolution Task. $\dagger$ denotes that we reproduce the results by retraining.
Models DIV2K Urban100 Time (s/image) Parameters Flops (G) NFE
PSNR SSIM LPIPS FID PSNR SSIM LPIPS FID
Bicubic 28.17 0.774 0.31 108.2 23.17 0.658 0.34 134.3 - - - -
RealESRGAN+ (Wang et al., 2021b) 28.18 0.776 0.115 13.57 24.37 0.734 0.123 20.29 0.0327 16.7M 73.43G 1
CAL-GAN (Park et al., 2023) 28.96 0.790 0.108 13.22 25.33 0.762 0.117 18.47 0.0327 16.7M 73.43G 1
SeD (Li et al., 2024a) 29.27 0.803 0.094 12.51 25.93 0.780 0.107 16.39 0.0327 16.7M 73.43G 1
RCoT (Tang et al., 2024) 29.63 0.819 0.279 20.04 24.87 0.742 0.270 33.63 0.0164 6.74M 204.74G 1
PromptIR (Potlapalli et al., 2023) 30.73 0.844 0.262 16.75 26.71 0.804 0.209 26.41 0.0602 33.19M 11.91G 1
Restormer (Zamir et al., 2022) 30.73 0.844 0.261 16.65 26.74 0.804 0.209 26.11 0.0557 26.32M 10.84G 1
RCAN (Zhang et al., 2018b) 30.78 0.844 0.255 16.41 26.74 0.807 0.196 24.23 0.0863 15.4M 65.25G 1
IPT (Chen et al., 2021) 30.92 0.847 0.253 15.76 27.33 0.820 0.189 22.92 0.0379 115M 43.41G 1
SwinIR (Liu et al., 2021) 31.08 0.852 0.247 15.29 27.46 0.826 0.184 21.98 0.0384 11.8M 37.17G 1
HAT (Chen et al., 2023a) 31.22 0.855 0.245 15.33 27.97 0.836 0.173 20.49 0.0613 20.6M 102.4G 1
MambaIRv2 (Guo et al., 2024a) 31.24 0.858 0.246 15.94 28.07 0.838 0.171 20.83 0.0626 34.2M 82.3G 1
SR3 (Saharia et al., 2022b) 20.37 0.629 0.370 46.04 20.60 0.635 0.288 43.55 50.4 155.3M 155.2G 100
SRDiff (Li et al., 2022b) 28.62 0.789 0.130 16.93 25.14 0.758 0.138 20.65 4.5 13.2M 84.22G 100
Resshift (Yue et al., 2023) 29.38 0.805 0.130 15.41 25.68 0.766 0.122 20.19 - - - 15
DiffBIR (Lin et al., 2024) 25.65 0.702 0.187 26.45 23.49 0.676 0.204 37.68 4.65 1716.71M 189.47G 50
SUPIR (Yu et al., 2024) 27.19 0.728 0.192 24.94 23.26 0.669 0.182 33.38 7.64 4801.1M 214.43G 50
Table 3: Quantitative Results of Zero-shot Models on Super-resolution Task. Best performances are bolded.
Models ImageNet 1K CelebA-HQ test Time (s/image) Flops (G) NFE
PSNR SSIM LPIPS FID PSNR SSIM LPIPS FID
Bicubic 25.98 0.699 0.461 87.56 28.91 0.786 0.355 154.69 - - -
RealESRGAN (Wang et al., 2021b) 25.94 0.705 0.123 30.29 29.48 0.801 0.099 26.60 0.0327 73.43G 1
CAL-GAN (Park et al., 2023) 26.92 0.736 0.122 29.14 30.00 0.815 0.101 22.40 0.0327 73.43G 1
SeD (Li et al., 2024a) 27.01 0.741 0.109 27.10 29.62 0.807 0.089 18.96 0.0327 73.43G 1
RCOT (Tang et al., 2024) 27.36 0.755 0.297 57.42 30.28 0.825 0.229 75.13 0.0164 204.74G 1
PromptIR (Potlapalli et al., 2023) 28.35 0.787 0.274 50.13 30.99 0.842 0.224 62.79 0.0602 11.91G 1
Restormer (Zamir et al., 2022) 28.34 0.786 0.274 49.78 30.98 0.842 0.224 62.23 0.0557 10.84G 1
RCAN (Zhang et al., 2018b) 28.35 0.786 0.269 50.92 30.95 0.839 0.222 60.87 0.0863 65.25G 1
IPT (Chen et al., 2021) 28.39 0.787 0.268 50.26 30.98 0.839 0.220 59.47 0.0379 43.41G 1
SwinIR (Liu et al., 2021) 28.56 0.792 0.263 49.57 31.03 0.841 0.220 60.72 0.0384 37.17G 1
HAT (Chen et al., 2023a) 28.67 0.795 0.262 48.20 31.08 0.841 0.220 60.69 0.0613 102.4G 1
MambaIR (Guo et al., 2024b) 28.62 0.793 0.264 48.87 31.07 0.843 0.219 59.36 0.0626 82.30G 1
ILVR (Choi et al., 2021) 27.40 0.871 0.21 43.76 31.59 0.878 0.22 32.24 41.3 1113.75 250
SNIPS (Kawar et al., 2021) 24.31 0.684 0.21 124.54 27.34 0.675 0.27 105.19 31.4 - 1000
DDRM (Kawar et al., 2022a) 27.38 0.869 0.22 40.75 31.64 0.946 0.19 31.04 10.1 1113.75 20
DPS (Chung et al., 2023b) 25.88 0.814 0.15 39.24 26.95 0.878 0.18 29.97 141.2 1113.75 1000
DDNM (Wang et al., 2023g) 27.46 0.871 0.15 40.15 31.64 0.945 0.16 28.27 15.5 1113.75 100
GDP (Fei et al., 2023) 26.51 0.832 0.14 38.45 28.65 0.876 0.17 27.51 3.1 1113.76 20
Reference
DDRM
DDNM
DPS
GDP
DiffPIR
StableSR
Figure 12: Qualitative results of diffusion models on Image Super-resolution
5.2 Experimental Results
To demonstrate the superiority of different diffusion models, we provide an objective quality comparison for them on multiple tasks. Concretely, we select three commonly investigated IR tasks, including image super-resolution, image deblurring, and image inpainting. The evaluation metrics are composed of PSNR, SSIM (Wang et al., 2004), FID (Heusel et al., 2017), and LPIPS (Zhang et al., 2018a). To compare the computation cost and network complexity, we also measure the running time, parameters, and flops for diffusion model-based IR methods. The subjective comparisons of image deblurring, image inpainting, and real-world IR are shown in Fig. 12, Fig. 3, 4, 5, 6 and 7 of section D.6 of the Supplementary. The comparison of diffusion-based methods on image inpainting and severe composite degradation can be found in the section D.4, D.5 of the Supplementary.
Results on Image SR and Real Image Restoration/SR. The experimental settings are as: (i) For the comparison of supervised diffusion model (DM)-based super-resolution (SR) methods, we use full images from DIV2K(Agustsson and Timofte, 2017) and Urban100(Huang et al., 2015) as test datasets instead of their cropped versions. We retrain SR3 (Saharia et al., 2022b) and SRDiff (Li et al., 2022b) with DF2K (Agustsson and Timofte, 2017) dataset to keep fair comparison. Additionally, we introduce two widely recognized diffusion-based Real-world SR methods, DiffBIR (Lin et al., 2024) and SUPIR (Yu et al., 2024), which are used to check the generalization capability of diffusion-based real-world image SR methods on synthetic datasets with bicubic degradation. (ii) For the comparisons of zero-shot DM-based SR methods, we follow the experimental setups of existing works, i.e., DPS(Chung et al., 2023b) and DDNM(Wang et al., 2023g), to process ImageNet1K (Russakovsky et al., 2015a) and CelebA-HQ test (Karras et al., 2017b) for consistent comparison. (iii) For the comparison of real-world image restoration, we follow StableSR (Wang et al., 2023c) to select RealSR (Cai et al., 2019), DRealSR (Wei et al., 2020), and RealPhoto60 as test datasets with central cropping operation.
[…]
For zero-shot-based diffusion IR models, the quantitative comparisons are shown in Table 3. We test six open-source diffusion models for 4x super-resolution on two datasets: ImageNet 1K (Russakovsky et al., 2015a) and CelebaHQ test (Liu et al., 2015). From the table, we can observe that DDRM (Kawar et al., 2022a) and DDNM (Wang et al., 2023g) performed well on various metrics, followed by ILVR (Choi et al., 2021). This is because DDRM and DDNM consider data consistency with the low-resolution image from a decomposition perspective, while ILVR only ensures low-frequency consistency. DPS (Chung et al., 2023b) and GDP (Fei et al., 2023) focus more on the perceptual quality of generated images. GDP (Fei et al., 2023) performs well in perceptual metrics on both datasets because it uses an empirical formula rather than Tweedie’s formula for posterior estimation, resulting in some performance improvement relative to DPS. GDP also exhibits the fastest image generation speed, with the reverse sampling steps of 20. The reason for that is it does not need the complex SVD decomposition and computation in DDRM, thereby accelerating image generation. On the other hand, DPS performs corrections after every sampling step (around 1000 steps), making it unable to utilize the DDIM-based sampling method for acceleration. Therefore, generating a single image with DPS takes approximately 141.2 seconds.
By comparing GAN-based IR methods and Diffusion-based methods, we can find that: (i) Diffusion-based IR methods do not always outperform GAN-based IR methods for synthetic degradation. As shown in Table 2 and 3, SeD (Li et al., 2024a) consistently achieve optimal reference-based perceptual quality metrics, including LPIPS and FID, which reveals the potential of Gan-based IR methods on synthetic degradation since they also have the advantage on efficiency; (ii) GAN-based IR methods and Diffusion-based IR methods are susceptible to poor pixel-wise consistency in general (i.e., PSNR, and SSIM) since they are generation-based models. But, we can find some zero-shot DM-based methods in Table 3, such as DDRM (Kawar et al., 2022a) and  (Wang et al., 2023g) achieve great pixel-wise consistency. The reason might be that decomposition-based methods can accurately obtain the inverse matrix for bicubic downsampling. Moreover, the utilized diffusion model for zero-shot IR has been pre-trained on face datasets; (iii) Diffusion-based RealSR methods outperform GAN-based SR methods by a large margin across all NR-IQA metrics for real-world degradation, as shown in Table 4, This could be attributed to the fact that these diffusion-based methods are built upon pre-trained diffusion models (e.g., Stable Diffusion (Wang et al., 2023c)) trained on large-scale natural image datasets. These models inherently possess strong priors that closely align with the real-world image distribution, enabling the super-resolution models fine-tuned on them to generate more realistic restoration results. Nevertheless, due to the nature of diffusion models, where images are sampled progressively from a Gaussian noise distribution, the stochastic generation process is difficult to constrain at the pixel level. This leads to inferior performance on objective metrics such as PSNR and SSIM compared to GAN-based methods, which can better preserve pixel-wise fidelity.
We also select seven state-of-the-art typical deterministic IR methods with various architectures for comparison, including transformer (Dosovitskiy et al., 2020), mamba (Guo et al., 2024b) and CNNs (Zhang et al., 2018b) in Tables 2 and 3. These methods achieve superior performance on objective metrics such as PSNR and SSIM compared to GAN-based and diffusion-based methods. However, this strict pixel-wise fidelity constraint comes at the cost of perceptual quality, leading to inferior performance on subjective evaluation metrics compared to GAN-based and diffusion-based approaches.
Results on Image Deblurring. We also evaluate five zero-shot DM-based IR methods on the Gaussian deblurring task using the ImageNet 1K (Russakovsky et al., 2015a) and CelebaHQ test (Liu et al., 2015) datasets. The experimental results are shown in Table 8 of the Supplementary. We can find that DiffPIR (Zhu et al., 2023) and Dirac-DO (Fabian et al., 2023) achieve competitive performance on PSNR and SSIM, with an average improvement of 1.0 dB to 1.4 dB over DDRM (Kawar et al., 2022a) and DDNM (Wang et al., 2023g). Moreover, Dirac-PO (Fabian et al., 2023) and DiffPIR (Zhu et al., 2023) show superior performance on perceptual metrics. DPS (Chung et al., 2023b) performs well on perceptual metrics, including LIPS and FID, but has a long generation time for each image. DiffPIR utilizes a plug-and-play sampling method and merges the DDIM sampling strategy to ensure both the fidelity and realism of generated images while speeding up the sampling process. Diracdiffion includes perception-optimized (PO) and distortion-optimized (DO) models, and employs an incremental reconstruction and early stopping method to achieve a perception-distortion trade-off. As a result, both models perform exceptionally well on distortion and perception metrics. Among all the models used, DDRM has the shortest sampling time, averaging under 10 seconds per image, as it only uses 20 sampling steps. In terms of model parameters, all zero-shot DM-based IR methods use a pre-trained model with 552.8M parameters for the ImageNet dataset and 126M parameters for the CelebA dataset.
[…]
Models RealSR DRealSR RealPhoto60 NFE
PSNR SSIM LPIPS MUSIQ ClipIQA ManIQA PSNR SSIM LPIPS MUSIQ ClipIQA ManIQA MUSIQ ClipIQA ManIQA -
RealESRGAN (Wang et al., 2021b) 25.69 0.762 0.271 60.37 0.448 0.548 28.64 0.805 0.282 54.27 0.451 0.490 50.95 0.599 0.488 1
CAL-GAN (Park et al., 2023) 26.38 0.754 0.272 56.07 0.423 0.505 28.91 0.799 0.289 50.00 0.423 0.460 51.52 0.646 0.522 1
SeD (Li et al., 2024a) 25.56 0.754 0.271 61.03 0.540 0.546 28.35 0.794 0.285 55.75 0.530 0.500 57.14 0.690 0.519 1
CDC (Wei et al., 2020) 27.81 0.787 0.329 42.71 0.299 0.321 32.76 0.877 0.291 40.27 0.356 0.283 29.52 0.369 0.268 1
SwinIR (Liu et al., 2021) 26.87 0.800 0.291 53.73 0.306 0.349 29.06 0.834 0.329 45.42 0.330 0.307 30.82 0.332 0.244 1
MambaIR (Guo et al., 2024b) 26.72 0.766 0.265 53.38 0.396 0.299 29.01 0.805 0.302 45.35 0.386 0.288 31.46 0.351 0.253 1
StableSR (Wang et al., 2023c) 24.70 0.707 0.300 65.88 0.623 0.428 28.15 0.752 0.328 58.51 0.636 0.387 62.24 0.651 0.361 200
DiffBIR (Lin et al., 2024) 25.07 0.640 0.395 64.85 0.638 0.615 26.53 0.630 0.495 61.19 0.634 0.592 64.45 0.689 0.418 50
PASD (Yang et al., 2024b) 24.29 0.663 0.343 68.69 0.659 0.649 27.00 0.708 0.393 61.81 0.677 0.585 63.28 0.712 0.477 20
DiffIR (Xia et al., 2023) 26.52 0.767 0.256 58.60 0.416 0.351 29.18 0.808 0.271 50.54 0.426 0.311 49.41 0.621 0.353 4
ResShift (Yue et al., 2023) 25.69 0.736 0.328 56.89 0.536 0.347 27.12 0.741 0.387 51.24 0.539 0.325 36.41 0.566 0.289 15
SUPIR (Yu et al., 2024) 23.47 0.537 0.518 67.06 0.663 0.641 23.76 0.563 0.659 65.56 0.592 0.581 67.04 0.746 0.515 50
6 Challenges and Future Directions
Although recent researches have achieved remarkable progress on diffusion model-based IR, there are still some challenges to extending them to practical applications since their limited robustness, model complexity, running efficiency, interactability, personalization, and restoration capability. To further improve the development of image restoration, we summarize the primary challenges and propose the potential directions for solving them in this section.
6.1 Sampling Efficiency
It is noteworthy that sampling efficiency is one typical challenge for the diffusion model, where the few sampling steps will cause limited generation fidelity. This inherent problem of the diffusion model damages the training and inference speed of image restoration. As shown in Table 2, SR3 (Saharia et al., 2022b) needs to take about 50 seconds to restore one image with the size of $224\times 224$, which is largely slower than the existing IR methods. previous works on diffusion models have attempted to improve the sampling efficiency from four perspectives: 1) modeling the diffusion process with a non-Markov Chain, such as DDIM (Song et al., 2020a). 2) designing efficient ODE solvers, e.g., DPM-solver. 3) leveraging the knowledge distillation to reduce sampling steps (Zhou and Tulsiani, 2023; Salimans and Ho, 2022), and 4) introducing the cross-modality priors with condition mechanism (Abu-Hussein et al., 2022; Wang et al., 2023c; Liu et al., 2023b). Under the above progresses, the sampling steps for the diffusion model are largely reduced to $10\sim 20$ steps, which also serves for more fast image restoration. In particular, DDRM (Kawar et al., 2022a) reduces the inference speed to 8 seconds for one $224\times 224$ of the image with the sampling strategy of DDIM (Song et al., 2020a).
Despite that, the above strategies are not specific to image restoration tasks. Differently, considering the low-quality image in IR contains abundant structural and textual information, some works (Luo et al., 2023b; Welker et al., 2022; Luo et al., 2023c; Zhao et al., 2023a) achieve image restoration by sampling from the low-quality image instead of pure noise, which obviates the extra sampling steps in the original DDPM. Among them, PartDiff (Zhao et al., 2023a) utilizes an intermediate latent stage containing Low-resolution information as a proxy for the start of the inference process. Another work (Ma et al., 2023) tries to accelerate the sampling speed of SR models by exploring the optimal boundary conditions for solving diffusion ODEs.
[…]
6.2 Model Compression
Model size is also a significant factor impacting the computational cost, limiting the real-time application of diffusion model-based image restoration (IR), such as mobile devices. In particular, DDPM (Ho et al., 2020) and SR3 (Saharia et al., 2022b) have 113.7M and 155.3M parameters, respectively, substantially exceeding previous CNN-based (Zhang et al., 2018b; Jiang et al., 2021a; Li et al., 2020) or Transformer-based IR backbones (Liu et al., 2021, 2022b; Wang et al., 2022e; Li et al., 2022a). To mitigate this, diffusion model compression is a potential but under-explored research direction toward efficient IR. Model compression (Li et al., 2023e), with the goal of reducing computational cost while maintaining task performance, has achieved great breakthroughs from four perspectives: including 1) model pruning, with the aim of removing the unimportant parameters by estimating the importance score of each parameter, 2) model quantization targets for reducing the bit-depth of the floating-point parameters for storage or computing, 3) knowledge distillation is proposed to transfer the knowledge from the complex teacher model to the simple and efficient student model, and 4) low-rank decomposition is devoted to decomposing the parameter tensor into multiple low-rank tensors. On basis of these, some works have taken a step forward to investigate the model compression of the diffusion model. Kim et al.(Kim et al., 2023a) introduce the block-removed knowledge distillation for the diffusion model, which constructs the student model by removing some residual and attention blocks from the U-Net architecture. Fang et al.(Fang et al., 2023) pinpoint that not all diffusion steps contribute to the generation process, and then, exploit partial diffusion steps to estimate the important and unimportant weights for parameter pruning with Taylor expansion. Additionally, there are pioneering works (Shang et al., 2023b; Wang et al., 2023a) that take the model quantization for the diffusion model to accelerate the sampling process. Although such progress, few works explore how to design the model compression for diffusion model-based IR, which is expected to be developed for real-time application.
6.3 Distortion Simulation and Estimation
Real-world/blind IR is a challenging yet significant task, with the goal of addressing the unknown and intricate degradations encountered in the real world. Unlike synthetic degradation, where the distortion is predefined and paired training samples are available, collecting the paired real-world distorted/clean pairs is non-trivial, thereby preventing the training with supervised learning. To address this limitation, unsupervised learning has been introduced to leverage unpaired real-world distorted/clean images. However, this learning paradigm usually yields an unsatisfied texture consistency between the restored image and the low-quality image. In contrast, distortion simulation serves as another efficient strategy to maintain supervised learning by simulating real-world degradations. Typically, RealESGRAN (Wang et al., 2021b) and SR3+ (Sahak et al., 2023) are the representative work to explore the hand-crafted second-order degradations for Real-world IR. Notwithstanding, the hand-crafted distortion simulation is hard to cover all the degradations in the real world. To mitigate this, some works are inspired by domain translation, and introduce the GAN/diffusion model to translate the synthetic distorted image to the real-world one or translate the real-world distorted image to the synthetic one. The former intends to simulate the real-world training pairs for supervised learning (Yang et al., 2023b), while the latter aims to directly utilize the IR network trained with synthetic images (Wei et al., 2023).
Another crucial challenge for real-world/blind IR stems from distortion estimation, which involves explicitly/implicitly identifying the distortion types or levels. In this work, we summarize the utilization of distortion estimation from two perspectives: 1) distortion adaptive learning and 2) solving the inverse problem in IR. From the first perspective, a notable example is kernel prediction for blind IR (Chung et al., 2023a), where the estimated kernel/representation is used to guide the adaptation of the pre-trained IR model to the unknown degradation. Inspired by this, if we can estimate the distortion type or degree in an explicit/implicit manner, we can achieve the unified IR framework based on distortion adaptive learning. For the second perspective, as stated in Sec. 3.4, lots of zero-shot diffusion model-based IR methods are based on the modelling of linear reverse problem. This poses the requirements for the identifying of degradation mode, which is necessary for the consistency constrain in diffusion model. Therefore, most of them are devoted into the synthetic distortions since the distortion mode in real world are hard to be identify. One distortion estimation technique is urgently developed to extend the zero-shot diffusion model-based IR to the real-world application.
6.4 Distortion Invariant Learning
Recently, we have witnessed the fast evolution of diffusion model-based IR for specific degradation. However, it inevitably suffers from unsatisfied robustness when applied to unseen distortion types and degrees. This raises a foundational question: How to achieve consistent image restoration across diverse distortion types and levels? To achieve this, we propose one direction termed distortion invariant learning (DIL) (Li et al., 2023c), aiming to enable the IR model to be generalized to unknown and diverse degradations. The principle of DIL is to learn the representation that is invariant under various degradation modes and preserve enough structure and textual information for reconstruction.
Inspired by domain generalization (DG) (Zhou et al., 2022; Gulrajani and Lopez-Paz, 2020; Li et al., 2018a; Wang et al., 2022a), we can present some potential methods to achieve DIL for IR by regarding each distortion mode as one domain. In the DG field, there are three typical methods to learn the domain invariant feature, including domain alignment (Wang et al., 2021c; Li et al., 2018b; Shao et al., 2019), data augmentation (Xu et al., 2020; Mancini et al., 2020), and meta-learning (Wang et al., 2020a; Jia et al., 2020; Rahman et al., 2020; Albuquerque et al., 2019). In particular, domain alignment aims to align the representations of source and target domains through minimizing contrastive loss (Wang and Liu, 2021), Maximum Mean Discrepancy (MMD), or adversarial learning, etc. Data augmentation is exploited to extend the domain diversity and consistency, which enables the model to obtain the domain-invariant capability. Meta-learning aims to learn the domain invariant representation by aligning the gradient between different domains, which is from the optimization perspective. By regarding the distortion mode as one specific domain, we can obtain several strategies to achieve distortion invariant learning: 1) we can exploit the encoder-decoder architecture for IR and align the representations from different distorted images before the decoder. 2) the second strategy can learn the distortion invariant representation from distortion augmentation, i.e., simulating various distortions in the real world as much as possible. 3) optimize the empirical risk minimization in IR with meta-learning like (Li et al., 2023c).
For diffusion model based image restoration, the model is usually composed of two components, noise predictor and condition module. Therefore, we can achieve distortion invariant learning from two perspectives: 1) learning the distortion invariant noise predictor and 2) distortion invariant condition. Obviously, once we implement the distortion invariant condition, we can leave the noise predictor invariant in supervised IR or exploit the pre-trained diffusion model in zero-shot IR. Based on this, some pioneering works attempt to redesign the condition module to achieve the distortion invariant condition, e.g., DifFace (Yue and Loy, 2024) and DR2 (Wang et al., 2023i). Notably, the distortion invariant condition also relies on distortion invariant learning for better conditions, which still entails lots of effort for evolution in future work.
6.5 Framework Design
As the foundation of image restoration, how to design an effective and powerful IR framework is an ongoing and significant question. We can notice that the most recent diffusion model-based IR methods (Saharia et al., 2022b; Wang et al., 2023g; Kawar et al., 2022a; Chung et al., 2023b; Saharia et al., 2022a; Song et al., 2023; Fei et al., 2023; Li et al., 2022b) are designed based on the U-Net architecture from DDPM (Ho et al., 2020), and pursue better frameworks from three perspectives, i.e., the condition strategy (Chung et al., 2023b; Kawar et al., 2022a; Lugmayr et al., 2022; Zhang et al., 2022a; Luo et al., 2023b), generation space(Rombach et al., 2022; Luo et al., 2023c; Xia et al., 2023), noise predictor (Xia et al., 2023), respectively. The condition of the diffusion model in IR aims to introduce the structure and textual information from low-quality images. In the early work, SR3 directly select low-quality images as a condition with the concatenation. To improve the condition, some works (Niu et al., 2023; Guo et al., 2023a; Gao et al., 2023b; Jin et al., 2022; Shang et al., 2023a) improve the condition by designing pre-processing networks, such as feature extractors and pre-trained restoration networks. From the generation space, the framework is usually designed from four spaces, including image space, residual space, latent space, and frequency space. Among them, pixel-wise space can preserve more spatial structure and textual information, which can generate high-quality images (Niu et al., 2023; Shang et al., 2023a; Gao et al., 2023b) or residuals (Li et al., 2022b; Shang et al., 2023a; Liu et al., 2024b; Yue et al., 2023; Liu et al., 2023a), while owning higher computational cost and parameters. In contrast, latent space generation requires fewer computational costs. However, one well-designed encoder and decoder are crucial for latent space generation to make the trade-off between efficiency and fidelity. Frequency space has been widely applied in diffusion model-based image rest-oration (Moser et al., 2023a; Xu et al., 2023; Mirza et al., 2023; Wang et al., 2023b; Guan et al., 2023; Luo et al., 2023a), including wavelet transform, Fourier transform, etc. Compared with image-wise space, frequency space is more good at capturing the global contextual information, where the low-frequency refers to the structural information and the high-frequency represents the texture and style information. These methods can bring more detailed and high-quality image restoration. Following the DDPM (Ho et al., 2020), in most works, the noise predictor is based on the U-Net architecture. For supervised diffusion model-based IR, the modification of the noise predictor is usually achieved by increasing the number of residual blocks in the U-Net or adjusting the channel multipliers at different resolutions, such as SR3. Few works are devoted to investigating how to design brand-new architectures based on transformers for the noise predictor in diffusion model-based IR. Furthermore, how to design the unified and foundational architecture like the painter (Wang et al., 2023f) for diffusion model-based IR tasks is urgently required to be explored.
6.6 All-in-one Restoration
All-in-one image restoration aims to simultaneously address multiple diverse degradations within a single IR model, which has attracted lots of attention in recent years. The key challenge lies in enabling diffusion-based IR models to adapt to various degradations while effectively leveraging shared restoration knowledge across different degradation types. The typical work is DA-CLIP (Luo et al., 2024), which introduces the degradation-aware CLIP to extract the degradation-aware embedding as the condition for diffusion models. In contrast, MPerceiver (Ai et al., 2024) extracts the holistic representation and multi-scale detail representation with the textual and visual branches of CLIP as the condition. Diff-Restore (Zhang et al., 2024) also leverages the CLIP model to extract the degradation and semantic embedding but adopts a distinct conditioning strategy. For unknown hybrid degradation, AutoDIR (Jiang et al., 2024b) introduce the quality assessment model to identify which degradation should be removed and utilizes multiple steps for restoration. UIR-LoRA (Hu et al., 2022) leverages low-rank adaptation (LoRA) to encode task embeddings efficiently. Then it introduces a dynamic weighting strategy based on CLIP similarity between image features and predefined distortion text embeddings. In contrast, GenDeg (Rajagopalan et al., 2024) utilizes diffusion models to generate diverse distortions from clean images, expanding training data for all-in-one diffusion-based IR models, which further improves the performance of multiple IR tasks. Although the above works have achieved impressive progress in diffusion-based all-in-one image restoration, a new all-in-one diffusion paradigm needs to be developed based on efficient architecture and new multi-task learning paradigm like Mixtures-of-Experts (MoE).
6.7 Restoration Systems with LLM/MLLM-based Agents.
Recent studies have proposed agent-based image restoration frameworks to enhance system intelligence, flexibility, and human interactivity. Leveraging the powerful understanding and reasoning capabilities of large language models (LLMs) and large multimodal models (LMMs), these approaches introduce language-driven agents into the restoration pipeline. For instance, AgenticIR (Zhu et al., 2024), hybridAgents Li et al. (2025) and RestoreAgent (Chen et al., 2024a) employ frozen LLMs or vision-language models to analyze input degradations and dynamically select traditional restoration tools (e.g., SwinIR (Liang et al., 2021), FBCNN (Jiang et al., 2021b), Restormer (Zamir et al., 2022)). Building upon this foundation, recent works such as MAIR (Jiang et al., 2025) and 4K-Agent (Yusheng et al., 2025) integrate powerful diffusion-based models—such as DiffBIR (Lin et al., 2024), IR-SDE (Luo et al., 2023b), and DPS (Chung et al., 2023b)—into the agent’s toolset. In particular, 4K-Agent employs a frozen LLM/MLLM to perform multi-round reasoning and sequential tool invocation, treating diffusion models as central, not auxiliary, components. This design achieves superior performance, especially under compound, real-world degradations. These developments reveal a clear trajectory in image restoration: from static, task-specific models to intelligent agentic systems, and further to fully integrated multimodal frameworks that emphasize interpretability, generalization, and controllability. As such, the development of LLM/MLLM-based restoration age-nts is emerging as a highly promising direction in the field, where the diffusion-based IR models are expected to serve as core generative priors and cross-modality generation tools in this system.
6.8 Vision-Language Prompting for Diffusion-Based Restoration.
The integration of vision-language models (VLMs) into restoration workflows enables high-level semantic guidance for diffusion models. PASD (Yang et al., 2024b) and Upscale-A-Video (Zhou et al., 2024) utilize frozen VLMs like BLIP-2 (Li et al., 2023a) and LLaVA (Liu et al., 2024a) to extract text prompts from degraded images, which are then used to condition diffusion upsamplers. Diff-Restorer (Zhang et al., 2024) takes this further by injecting semantic and degradation embeddings into the diffusion process. However, since standard VLMs may be misled by visual noise, recent methods like SeeSR (Wu et al., 2024b) and SUPIR (Yu et al., 2024) adapt VLMs to be degradation-aware by retraining (e.g., RAM-based tagging in SeeSR) or introducing lightweight artifact removers. These methods demonstrate the semantic understanding with VLMs, and the VLM-based condition strategies will empower high-fidelity restoration, which is a potential direction for diffusion-based IR.
6.9 Human-in-the-Loop Optimization and Preference Alignment.
Most existing image restoration (IR) models are optimized using L1, perceptual, or similar handcrafted losses, which may not accurately reflect human preferences or perceptual quality. To bridge this gap, recent works have begun to explore diffusion-based generation with human feedback. For instance, PrefPaint (Liu et al., 2024c) learns a reward model from human-anno-tated preference data and fine-tunes a pre-trained diffusion backbone via reinforcement learning. InstructRestore (Liu et al., 2025) enables region-specific, instruction-driven restoration by allowing users to specify both spatial regions and restoration strength using natural language. DSPO (Cai et al., 2025) introduces patch-level semantic preference optimization for real-world super-resolution, aligning local instance appearance with human feedback. In the future, the human-in-the-loop optimization is expected to open up an exciting direction, such as reinforcement learning with human feedback (RLHF), synthetic preference data generation via diffusion sampling, and fine-grained/region-aware diffusion-based IR restoration under natural-language control.
7 Conclusions
This work presents a comprehensive review of recent popular diffusion models for IR, excavating their substantial generative capability to enhance structure and texture restoration. Initially, we illustrate the definition and evolution of the diffusion model. Subsequently, we provide a systematic categorization of existing works from the perspectives of training strategy and degradation scenario. Concretely, we group existing works into three prominent flows: supervised DM-based IR, zero-shot DM-based IR, and DM-based blind/real-world IR. For each flow, we provide the fine-grained taxonomy based on the techniques and delicately describe their advantages and disadvantages. For evaluation, we summarize the commonly used datasets and evaluation metrics for DM-based IR. And we compare the open-sourced SOTA methods with distortion and perceptual metrics on three typical tasks, including image SR, deblurring, and inpainting. To overcome the potential challenges in DM-based IR, we highlight nine potential directions expected to be explored in the future.
Acknowledgements
This work was partly supported by the NSFC under Grant No. 623B2098, 62371434, 62021001, and the China Postdoctoral Science Foundation-Anhui Joint Support Program under Grant Number 2024T017AH.
We would like to appreciate Prof. Chen Chang Loy sincerely for his valuable suggestions and help on this paper, which greatly improved the quality of our work.
Data Availability Statements
All datasets and codes used in this work are publicly available and are cited in this work. We have summarized them in https://drive.google.com/file/d/1z_k1O6yFznDQ1OuyJuazK-YOauA8XHYt/view?usp=drive_link, which can be utilized to reproduce all experimental results of this work.